Grade 11Physical Chemistry

One mole of an ideal gas diatomic absorbs 1831.4 Joule of heat and does 1000 j of work at 27c .what is change in temperature

Sai Vamsi
9 Years agoGrade 11

1 Answer

Khimraj
9 Years ago
dU = dQ +dW. So RdT = 1831.4-1000. and then dT = 100°C.Hope it clears. If u have doubts then please clearify. And if u like my answer then please approve answer